OpenStudy (anonymous):

Evaluate the following expressions. Your answers must be exact and in simplest form. lne^-11=? e^(ln4)=? e^(lnsqrt(3))=? ln(1/(e^2))=? i would like to know the steps please:)

OpenStudy (anonymous):

\[ln(e^x)=x\] so \[ln(e^{-11})=-11\]

OpenStudy (anonymous):

similarly \[e^{ln(x)}=x\] so \[e^{ln(4)}=4\]

OpenStudy (anonymous):

and \[e^{ln(\sqrt{3})}=\sqrt{3}\]

OpenStudy (anonymous):

for the last one, \[ln(\frac{1}{x})=-ln(x)\] so \[ln(\frac{1}{e^2})=-ln(e^2)=-2\]

OpenStudy (anonymous):

btw \[ln(\frac{a}{b})=ln(a)-ln(b)\]so \[ln(\frac{1}{b})=ln(1)-ln(b)=-ln(b)\]because \[ln(1)=0\]

OpenStudy (anonymous):

thank you !:)

OpenStudy (anonymous):

welcome. hope it is clear!

OpenStudy (anonymous):

oh yea def going in my notes lol :D

OpenStudy (anonymous):

notes? ;)

OpenStudy (anonymous):

yea so i can remember them and cram them in my head hahaha

OpenStudy (anonymous):

well its easy just gotta remember haha